OpenStudy (anonymous):

How do I get: 12(x-4)-(x+2)=(x+2)(x-4) into quadratic standard form? I have the answer for what the standard form is for this, I just have no clue how they got it. Book is fail.

OpenStudy (anonymous):

standard form is \[x ^{2}-13x+42=0\] if anyones wondering

OpenStudy (chaise):

12(x-4)-(x+2)=(x+2)(x-4) Let's start by expanding the brackets? (12x-48)-(x+2)=(x+2)(x-4) 11x-50=(x+2)(x-4) Let's foil out the right hand side? 11x-50=x^2-2x-8 Minus 11x both sides, plus 50. both sides. x^2-13x+42=0 Too easy?

OpenStudy (anonymous):

thank you for explaining. Book just threw it at me without any explination